OpenStudy (anonymous):

What is the distance between the points (-5, 2) and (4, 9)? >>

OpenStudy (anonymous):

\[d=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\]

OpenStudy (anonymous):

A)

OpenStudy (anonymous):

B)

OpenStudy (anonymous):

11.4

OpenStudy (anonymous):

C)

OpenStudy (anonymous):

\[\sqrt{130}\]

OpenStudy (anonymous):

C

OpenStudy (anonymous):

D)

OpenStudy (anonymous):

\[\sqrt{(-5-4)^2+(2-9)^2}=\sqrt{81+49}=\sqrt{130}\]

OpenStudy (anonymous):

EASY